OpenStudy (gamenerd123):

Initially, a store has 35 train sets. The store sells 3 train sets each month. What is a explicit rule for the number of train sets the store has after n months? An=?

OpenStudy (welshfella):

this is an arithmetic sequence a1 (first term) = 35 d (common difference) = -3

OpenStudy (gamenerd123):

So what will that mean?

OpenStudy (welshfella):

an = a1 + d(n - 1)

OpenStudy (gamenerd123):

Oh.

OpenStudy (welshfella):

- well just plug those values into the general formula (in my last post)

OpenStudy (gamenerd123):

Okay

OpenStudy (mathmale):

Note how welshfella has come up with an equation that enables you to calculate how many train sets are still available after n months? What is the initial value of that function? What is the common difference? welshfella knows what he's doing. Pls share your work.

OpenStudy (mathmale):

\[nth ~term ~ is ~a _{n}\]

OpenStudy (gamenerd123):

So what I come up with will be the answer

OpenStudy (mathmale):

\[First ~ is ~a _{1}\]

OpenStudy (mathmale):

common difference, d, is what? What does n represent?

OpenStudy (gamenerd123):

d(n-1)?

OpenStudy (mathmale):

as welshfella has said, plug in the knowns to obtain a formula. d(n-1) is part of your formula. What's the whole formula? The purpose of this formula is to calculate the nth term of your sequence.

OpenStudy (gamenerd123):

Well I'm very familiar of how to finish it tbh

OpenStudy (mathmale):

That statement seems to say, "I'm sure I can finish it." Is that what you meant to say?

OpenStudy (gamenerd123):

No, I'm not sure how to finish it, sorry

OpenStudy (mathmale):

From welshfella: an = a1 + d(n - 1) where the first term is 35, the common diff is d, the index of the answer you want is n, and an is the nth term. Start substituting these knowns (first term and common difference) into the equation.

OpenStudy (gamenerd123):

Ahh okay

OpenStudy (gamenerd123):

3(n-35)?

OpenStudy (mathmale):

Please type out the entire formula. Start it with \[a _{n}=\]

OpenStudy (mathmale):

Next, write in the value of the first term (35).

OpenStudy (mathmale):

Next, write in d(n-1), but before you do that, replace "d" with the "common difference."

OpenStudy (gamenerd123):

The common difference is 35, right?

OpenStudy (mathmale):

\[a _{n}=35+????\]

OpenStudy (mathmale):

No, the first term is 35. The store starts out with 35 train sets in stock and sells 3 sets each month. What is the common difference? Is the # of train sets increasing or decreasing?

OpenStudy (gamenerd123):

Well it seems to be decreasing when they sell 3 sets each month

OpenStudy (mathmale):

If increasing, then the common diff is +, but if decreasing, then the com. diff. is - .

OpenStudy (mathmale):

So, what is the common difference?

OpenStudy (mathmale):

d=?

OpenStudy (gamenerd123):

3?

OpenStudy (mathmale):

Is the # of train sets in stock in the store incr or decr?

OpenStudy (gamenerd123):

Its decreasing

OpenStudy (mathmale):

Then the common diff must be negative. What is the common diff?

OpenStudy (gamenerd123):

Oh, so must I subtract 35 and 3?

OpenStudy (mathmale):

No. You begin with \[a _{n}=35+d(n-1)\] and replace "d" with "-3." Do that now, please. Keep in mind what these numbers stand for. 35 is the starting count of the number of sets on hand. -3 is the common difference, d. n is the count: 1, 2, 3, and so on.

OpenStudy (gamenerd123):

Oh, I'm sorry for giving a hard time, sorry

OpenStudy (mathmale):

You're learning, so I'm happy.

OpenStudy (gamenerd123):

So how would I write that equation?

OpenStudy (mathmale):

\[a _{n}=35+(-3)(n-1)\]

OpenStudy (mathmale):

Please simplify: " +(-3) "

OpenStudy (mathmale):

(+)(-)=?

OpenStudy (gamenerd123):

lol I have no clue of how to simplify

OpenStudy (mathmale):

(+)(-)=(-) So you end up with \[a _{n}=35-3(n-1).\]

OpenStudy (mathmale):

This is the answer you wanted. Please, review this entire discussion and ask any necessary questions that might help you arrive at a complete understanding.

OpenStudy (gamenerd123):

Ahh okay I get it now

OpenStudy (gamenerd123):

I wrote it all down in notes

OpenStudy (mathmale):

This formula is "arithmetic sequence," that is, a formula for the nth term of an arith sequence when the starting value (a_1), the common diff (d) are known. n is just a counter: it begins with 1, moves on to 2, 3, 4, .... and so on.

OpenStudy (gamenerd123):

Okay, I'll write that down, the post another subject, one sec..

OpenStudy (mathmale):

\[a _{n}=?\]

OpenStudy (mathmale):

can you now define what that means?

OpenStudy (gamenerd123):

As for this problem?

OpenStudy (mathmale):

Yes. What does that a_n represent? answer in words only.

OpenStudy (gamenerd123):

It represents N=1

OpenStudy (mathmale):

It represents "the nth term of an arithmetic sequence."

OpenStudy (mathmale):

a_1 is the first term, a_2 is the second term, and so on. what does "d" stand for?

OpenStudy (gamenerd123):

From what I wrote, D stands for the Difference, am I right?

OpenStudy (mathmale):

common difference. Yes.

OpenStudy (mathmale):

What does "n" stand for?

OpenStudy (gamenerd123):

Oh, common difference lol sorry

OpenStudy (gamenerd123):

N is the number count for 1,2,3,4,5, and so forth

OpenStudy (mathmale):

Yes. we call it an "index." for example, I might say, "Hey, Nerd, give me the 4th member of this arith sequence." n would be 4. What does a_1 stand for?

OpenStudy (gamenerd123):

It represents the first term of the Arithmetic sequence

OpenStudy (mathmale):

Right. And in this particular problem (case), what is the value of a_1?

OpenStudy (gamenerd123):

3?

OpenStudy (mathmale):

Again, what does a_1 represent? '3' is something else.

OpenStudy (mathmale):

You could, of course, go back and review this discussion if necessary, to review what a_1 represents.

OpenStudy (gamenerd123):

Yes I just did, it represents one, right? I don't believe I wrote that down in my notes

OpenStudy (mathmale):

a_1 is the starting value of this arith sequence, and is 35.

OpenStudy (gamenerd123):

Which is the first term

OpenStudy (mathmale):

Yes, it's the first term.

OpenStudy (gamenerd123):

Oh

OpenStudy (gamenerd123):

I'm gonna write that down right now

OpenStudy (mathmale):

\[a _{n}=a _{1}-3(n-1)\]

OpenStudy (mathmale):

is the answer to this particular question. Hey, Nerd, what's the 7th term of this sequence?

OpenStudy (mathmale):

Let n=7 in this formula\[a _{n}=a _{1}-3(n-1)\]

OpenStudy (gamenerd123):

Ah okay

OpenStudy (mathmale):

7-1=?

OpenStudy (gamenerd123):

6

OpenStudy (gamenerd123):

Shall I post my next problem?

OpenStudy (mathmale):

Let's go thru that: a_7 = 35 - 3(6), or 35-? = ?

OpenStudy (mathmale):

Therefore, a_7 = ??

OpenStudy (mathmale):

a_7 = 35 - ??

OpenStudy (gamenerd123):

a_7=35

OpenStudy (gamenerd123):

3(6)

OpenStudy (mathmale):

but -3(6) = -18. So, a_7 = 35 - 18 = ??

OpenStudy (mathmale):

35 -18 ------ ?

OpenStudy (gamenerd123):

17

OpenStudy (mathmale):

yes, and so a_7 = 35 - 17 = ?

OpenStudy (gamenerd123):

18

OpenStudy (mathmale):

a_7 represents the number of remaining train sets after 7 weeks of sales of the trains.

OpenStudy (mathmale):

Yes. The store has 18 sets left after 7 weeks. Yes.

OpenStudy (gamenerd123):

That makes better sence lol

OpenStudy (mathmale):

Formulas, formulas. gotta learn to be comfy with them.

OpenStudy (mathmale):

algebraic representation, algebraic rep, gotta be comfy.

OpenStudy (gamenerd123):

Well its a little difficult cuz I hate Math XD

OpenStudy (mathmale):

Come on, you like math and you're liking it better day by day. ;)

OpenStudy (mathmale):

Altho' I cannot help you with your remaining questions right now, I'll be back on sev eral times later today. You could either post your next question now and wait for another helper to respond, or you could look for me on OpenStudy later in the day.

OpenStudy (gamenerd123):

lol okay, thank you so much